OpenStudy (anonymous):

one number is 4 less than 3 times a second number. If 3 more than two times the firt number is decreased by 2 times the second number, the result is 11. Use the substitution method. what is the first number?

OpenStudy (anonymous):

first one-> x second one-> y now from first sentence x=3y-4 3+2x -2y=11 3+2(3y-4) -2y=11 3+6y-8-2y=11 4y-5=11 4y=11+5=16 y=4 hence x=3x4-4 = 12-4 = 8 x=8, y=4 understood?? I hope you did and its correct